Integration by Partial Fractions


∫dy/(5+4cosy)


1
Expert's answer
2021-05-07T09:12:44-0400

"Given \ that \ \int(\frac{1}{5+4cosy})dy\\
Substituting \ cos(y)= \frac{1-tan^2{\frac{y}{2}}}{1+tan^2{\frac{y}{2}}}\\
we \ get \int(\frac{1}{5+4cosy})dy=\int(\frac{1}{5+4(\frac{1-tan^2{\frac{y}{2}}}{1+tan^2{\frac{y}{2}}})})dy\\
=\int(\frac{1+tan^2{\frac{y}{2}}}{5(1+tan^2{\frac{y}{2}})+4({1-tan^2{\frac{y}{2}}})})dy\\
=\int(\frac{1+tan^2{\frac{y}{2}}}{5+5tan^2{\frac{y}{2}}+4-4tan^2{\frac{y}{2}}})dy
=\int(\frac{1+tan^2{\frac{y}{2}}}{9+tan^2{\frac{y}{2}}})dy\\
Substituting \ u=tan{\frac{y}{2}} \& \ (1+tan^2{\frac{y}{2}})dy=2du\\
we \ get \int(\frac{2}{9+u^2})du=\frac{2}{3}tan^{-1}(\frac{u}{3})+C,\\
 \because \int(\frac{1}{a^2+x^2})dx=\frac{1}{a}tan^{-1}(\frac{x}{a})\\where \ C \ is \ a \ constant \ of \ integration.\\
Substituting \ back \ u=tan{\frac{y}{2}} \ , we \ get \\
\int(\frac{1}{5+4cosy})dy=\frac{2}{3}tan^{-1}(\frac{tan{\frac{y}{2}}}{3})+C\\
\therefore \int(\frac{1}{5+4cosy})dy=\frac{2}{3}tan^{-1}(\frac{1}{3}tan{\frac{y}{2}})+C"